OpenStudy (anonymous):

find the missing factor. b^2-3b-4=(b-4)(?)

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

i still dont get it?

OpenStudy (anonymous):

Simplifying b2 + 3b + -4 = 0 Reorder the terms: -4 + 3b + b2 = 0 Solving -4 + 3b + b2 = 0 Solving for variable 'b'. Factor a trinomial. (-4 + -1b)(1 + -1b) = 0 Subproblem 1 Set the factor '(-4 + -1b)' equal to zero and attempt to solve: Simplifying -4 + -1b = 0 Solving -4 + -1b = 0 Move all terms containing b to the left, all other terms to the right. Add '4' to each side of the equation. -4 + 4 + -1b = 0 + 4 Combine like terms: -4 + 4 = 0 0 + -1b = 0 + 4 -1b = 0 + 4 Combine like terms: 0 + 4 = 4 -1b = 4 Divide each side by '-1'. b = -4 Simplifying b = -4 Subproblem 2 Set the factor '(1 + -1b)' equal to zero and attempt to solve: Simplifying 1 + -1b = 0 Solving 1 + -1b = 0 Move all terms containing b to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -1b = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -1b = 0 + -1 -1b = 0 + -1 Combine like terms: 0 + -1 = -1 -1b = -1 Divide each side by '-1'. b = 1 Simplifying b = 1 Solution b = {-4, 1} does that help?

OpenStudy (anonymous):

b^2-3b-4=b^2-(4-1)b-4=b^2-4b+1b-4=b(b-4)+1(b-4)=(b-4)(b-1) [ 1x-4=-4 4-1=3 4x-1=-4 Write -3=-(4-1) ]

OpenStudy (anonymous):

+-b + or - sqrt (b^2 - (4)(a)(c) all divided by 2(a) -(-3) plus or minus sqrt (-3)^2 - (4)(1)(-4) divided by -2 3 +/_ sqrt 9+16 divided by -2 3 +/- sqrt 25 divided by -2 3 + 5/-2 = -4 3 - 5/-2 = 1 your x values are 1 and -4.

OpenStudy (anonymous):

sorry the first part should be -b not +-b typo